OpenStudy (anonymous):

3x + 2y = 5 5x − 3y = 2

OpenStudy (shamil98):

3x + 2y = 5 5x - 3y = 2 3(3x+2y =5) 2(5x-3y=2) 9x + 6y = 15 10x - 6y = 4 19x = 19 x= 1 3(1) + 2y = 5 2y = 2 y= 1

OpenStudy (anonymous):

3x+2y=5 5x-3y=2 Soling for y -3y=2-5x y=5/3x-2/3 Substitute the value of y into the first the equation 3x+2(5/3x-2/3)=5 3x+3.33x-1.33=5 6.33x=6.33

OpenStudy (anonymous):

x=1

OpenStudy (shamil98):

I've answered it. -_-